# Robust Federated Learning via Byzantine Filtering over Encrypted Updates

### v1

@@ Abstract
Federated Learning (FL) aims to train a collaborative model while preserving data privacy. However, the distributed nature of this approach still raises privacy and security issues, such as the exposure of sensitive data due to inference attacks and the influence of Byzantine behaviors on the trained model. In particular, achieving both secure aggregation and Byzantine resilience remains challenging, as existing solutions often address these aspects independently. In this work, we propose to address these challenges through a novel approach that combines homomorphic encryption for privacy-preserving aggregation with property-inference-inspired meta-classifiers for Byzantine filtering. First, following the property-inference attacks blueprint, we train a set of filtering meta-classifiers on labeled shadow updates, reproducing a diverse ensemble of Byzantine misbehaviors in FL, including backdoor, gradient-inversion, label-flipping and shuffling attacks. The outputs of these meta-classifiers are then used to cancel the Byzantine encrypted updates by reweighting. Second, we propose an automated method for selecting the optimal kernel and the dimensionality hyperparameters with respect to homomorphic inference, aggregation constraints and efficiency over the CKKS cryptosystem. Finally, we demonstrate through extensive experiments the effectiveness of our approach against Byzantine participants on the $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{GTSRB}$, and $\mathsf{ACSIncome}$ benchmarks. More precisely, our SVM filtering achieves accuracies between 9090% and 9494% for identifying Byzantine updates at the cost of marginal losses in model utility and encrypted inference runtimes ranging from $6$ to $24$ seconds and from $9$ to $26$ seconds for an overall aggregation.
@@ 1 Introduction
    We conduct extensive experiments to demonstrate the performance of our approach against Byzantines on the $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{GTSRB}$ and $\mathsf{ACSIncome}$ datasets, along with ResNet He et al. (2015), VGG Simonyan and Zisserman (2015), custom CNN and MLP model architectures  [^1] Our implementations are available at https://github.com/Akram275/FL_with_FHE_filtering. For instance, our filtering approach achieves F1-scores between 9090% and 9494%, at the cost of marginal losses in model utility and encrypted inference runtime ranging from $3$ to $7$ seconds for a global aggregation (cf. Section 9).
@@ 5 Encrypted Property Filtering Framework
FL with FHE property filtering. Before training, workers initialize a threshold CKKS scheme with parameter $\tau\leq B_{\text{coll}}$, each holding a secret key share $\mathsf{sk}_{i}$ and the public key pk. The evaluation and public keys $(\textsf{evk},\textsf{pk})$ are sent to the server with the collaborative decryption requiring at least $\tau$ shares. The threshold CKKS setup is detailed in Appendix A.2.
When the FL training begins ($t=0$), the server broadcasts the initial plaintext model. Afterwards, workers compute local updates $\mathsf{LocalUpdate(\cdot)}$, encrypt them under pk and transmit the ciphertexts to the server. Then, the server computes the encrypted difference $\llbracket\Delta^{t}_{i}\rrbracket=\llbracket\theta^{t,s}\rrbracket-\llbracket\theta^{t,s}_{i}\rrbracket$ in which $s$ is the identifier of the layer with the highest separability (c.f., Section 6). An FHE property inference process $\mathsf{Filter}_{P}(\cdot)$ on $\llbracket\Delta^{t}_{i}\rrbracket$ results in an encrypted filter value $\llbracket P_{i}^{t}\rrbracket$ in which $P_{i}^{t}\in[0,1]$, is the filter value for worker $i$. Each update is homomorphically scaled by its corresponding filter value, and the scaled updates are aggregated (via averaging) to form the encrypted non-normalized model $\llbracket\theta_{\text{scaled}}^{t+1}\rrbracket$. The server homomorphically computes $\llbracket S^{t}\rrbracket=\sum_{i=1}^{n}\frac{n_{i}}{\sum_{i=1}^{n}n_{i}}\llbracket P^{t}_{i}\rrbracket$ and its inverse $\llbracket S_{\text{inv}}^{t}\rrbracket$ to normalize the global model. Hence, producing $\llbracket\theta^{t+1}\rrbracket$, as outlined in Algorithm 4. Workers collaboratively decrypt the aggregated model using their private keys shares $\mathsf{sk}_{i}$. Concurrently, the server bootstraps the aggregated model’s ciphertext $s$ to compute the next round’s $\Delta^{t+1}_{i}$.
@@ 7.2 Optimizing the SVM Filtering
To showcase our approach, we consider four models for $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{ACSIncome}$ and $\mathsf{GTSRB}$ presented in Table III. Figure 4 depicts the enhanced separability brought by SPCA compared to PCA on backdoor update differences from the first convolution layer. Indeed, SPCA widens the separation along the first component between honest and backdoor update clusters, moving from an average distance of around $2$ to more than $200$. This separation directly impacts the SVM’s decision boundary and reduces the number of support vectors, thereby improving inference runtime.
@@ 7.3 Optimal filters via grid search
Since the evaluation of the sum over all support vectors $(\mathsf{SVs})$ during inference is highly parallelizable, the factor $\mathsf{P}$ adjusts the influence of $(n_{\mathsf{sv}})$ count on the overall score. This ensures that configurations with a higher number of support vectors are not unfairly penalized when parallel processing is achievable. The weight $\mathsf{P}$ ranges from $1$, indicating no parallelization capabilities, to $\frac{1}{n_{\mathsf{sv}}}$, representing maximum parallelization, in which each support vector is evaluated simultaneously, with intermediate values capturing varying parallelization capabilities.
@@ 8.2 Polynomial approximation of Heavyside
This process reduces the multiplicative complexity of the polynomial evaluation from $O(d)$ to $O(\sqrt{d})$, which leads to two key benefits. First, since multiplications are the most computationally expensive operation in CKKS (and FHE in general), reducing the number of multiplications directly enhances efficiency. Second, by lowering the overall multiplicative depth of our FHE circuit, we can use more efficient CKKS parameters, reducing the cost of all unitary FHE operations. For our Chebyshev approximation of $\mathsf{Sigmoid}$, we use $d=32$, which requires a homomorphic multiplicative depth of $8$.
@@
in which $\mathsf{B}\neq 0$ denotes a new boundary value. Negative values induce less tolerance, in which higher confidence is needed to achieve $P^{t}_{j}\geq 0.5$, while positive values reduce the tolerance of the filtering process. From an FHE perspective, this is a ciphertext-plaintext addition [^3] The $\mathsf{B}$ value is encoded as a degree $N$ polynomial and summed directly to the second polynomial term of the RLWE ciphertext format (c.f. Appendix A)., which is computationally negligible and does not increase the ciphertext’s noise, nor consumes FHE-computational levels. This flexibility allows the approach to be tailored to the specific property. For instance, when targeting gradient-ascent attackers, incorporating even a single Byzantine update into the aggregation can severely disrupt the achieved learning progress, making it a greater concern than temporarily down-weighting an honest worker for one round.
@@ 8.3 Model Normalization with Newton-Raphson
In Algorithm 2, each worker’s update $\theta^{t}_{i}$ is scaled by $\frac{n_{i}}{\sum_{k\in[n]}n_{k}}\llbracket P^{t}_{i}\rrbracket$. This approach requires a normalization step after computing the sum of these scaled updates. Indeed, the scaling factors do not necessarily sum to $1$ (unless all the property filters are equal to $1$). The encrypted sum of the factors $\llbracket S^{t}\rrbracket=\sum_{i\in[n]}\left(\frac{n_{i}}{\sum_{k\in[n]}n_{k}}\right)\llbracket P^{t}_{i}\rrbracket$ is in $[0,1]$. Hence, to readjust the aggregated model’s parameters to the correct scale, a division by $S^{t}$ is required at each iteration. Since the division is not directly supported in the CKKS encryption, we approximate the inverse of $S^{t}$ using the Newton-Raphson method.
@@
Figure 8 illustrates that $k=2$ yields a sufficiently accurate inverse approximation, with an expected approximation error of around $10^{-2}$ and a computational cost of $4$ FHE products. Thus, we consistently use $k=2$ for $\mathsf{EncryptedInverse}(\cdot)$ in our aggregation circuit.
@@ 9.1 Encrypted property-filtering
Table VI illustrates the practical applicability of our filtering method, showing that most FHE operations (excluding bootstrapping) are completed in $0.9$ to $\approx 15$ seconds. The SPCA projection, which processes raw, high-dimensional encrypted update vectors, is the most computationally expensive step. In contrast, the SVM hyperplane projection is much faster thanks to its lower input dimension. The Paterson-Stockmeyer optimization significantly speeds up the evaluation of the $32$-degree Chebyshev approximation of the $\mathsf{Sigmoid}$ function. As a result, the overall filter computation time remains under $9$ seconds for Backdoor, gradient-ascent, and label-shuffling SVMs. Label-flipping filtering requires a quadratic kernel SVM with a larger number of support vectors and parameters ($\mathsf{P2}$), leading to an increased computation time. Although bootstrapping is costly, it is performed on the server side and can overlap with worker-side operations, thereby reducing its impact on the FL workflow. Further details are given in Appendix B.
@@ A.1 The CKKS scheme
  • $\mathsf{ParamsGen}(\lambda,L,\text{precision})$: Generates the public parameters for the CKKS scheme based on the security parameter $\lambda$, the desired multiplicative depth $L$, and the target precision for approximate arithmetic. Outputs typically include: the polynomial ring degree $N$ (a power of 2, determining the vector length), the chain of ciphertext coefficient moduli $Q=q_{0}\cdot q_{1}\cdots q_{L}$ (where $Q_{l}=q_{0}\cdots q_{l}$ is the modulus at level $l$), an auxiliary modulus $P$ used for key switching (relinearization, rotation), the scaling factor $\Delta$ (often a power of 2, related to precision), and specifications for the secret key distribution $S$ and error distribution $\chi$. These parameters must ensure $\lambda$-bit security while supporting $L$ levels of multiplication and the desired precision.
@@
  • $\mathsf{RelinKeyGen}(\textsf{sk})$: Generates the relinearization key $\mathsf{rlk}$, which is necessary to reduce the size of ciphertexts after multiplication. It is effectively an encryption of $s^{2}$ under the secret key sk, typically constructed using the extended modulus $PQ$. A common form involves sampling $a^{\prime}\leftarrow R_{PQ}$, $e^{\prime}\leftarrow\chi$, yielding $\mathsf{rlk}=(b^{\prime},a^{\prime})=([-a^{\prime}s+e^{\prime}+Ps^{2}]_{PQ},a^{\prime})$.
  • $\mathsf{RotKeyGen}(\textsf{sk},k)$: Generates a rotation key $\mathsf{rotk}_{k}$ required for homomorphically rotating the encrypted vector by $k$ positions (where $k$ corresponds to a specific Galois element). This key effectively encrypts the modified secret key $s(X^{k})$ under the original secret key sk, typically using the modulus $PQ$. Keys must be generated for all desired rotation indices $k$.
@@ Appendix B Offline amortized cost of Bootstrapping
During the collaborative decryption phase, workers generate $t$-out-$t$ additive keys by (at least) $t$ active parties, using their respective $n$ shares of the initial $n$ additive secret keys (c.f. equation 4). Apart from revealing the active parties, the process of computing the $t$ partial decryptions does not require data exchanged other than what has already been exchanged during the setup. The last step is broadcasting the respective partial decryption by each of the $t$ active parties. A partially decrypted message consists of a degree $N$ polynomial with coefficients in $\mathbb{Z}_{Q}$. Thus, the size of data exchanged (for minimal $t$ active parties) is $t(t-1)N\log_{2}(Q)$ bits for a single ciphertext. Encrypting a model update requires $\left\lceil\frac{\mathsf{dim}(\theta)}{N/2}\right\rceil$ ciphertexts, which are 3 to 15 ciphertexts for our models (see table III). Table VII illustrates the size of the exchanged data during the collaborative decryption and the runtime necessary to generate the shares for several values of $\log(N)$, and different numbers of active parties ($t$).

### v2

@@ Abstract
Federated Learning (FL) aims to train a collaborative model while preserving data privacy. However, the distributed nature of this approach still raises privacy and security issues, such as the exposure of sensitive data due to inference attacks and the influence of Byzantine behaviors on the trained model. In particular, achieving both secure aggregation and Byzantine resilience remains challenging, as existing solutions often address these aspects independently. In this work, we propose to address these challenges through a novel approach that combines homomorphic encryption for privacy-preserving aggregation with property-inference-inspired meta-classifiers for Byzantine filtering. First, following the property-inference attacks blueprint, we train a set of filtering meta-classifiers on labeled shadow updates, reproducing a diverse ensemble of Byzantine misbehaviors in FL, including backdoor, gradient-inversion, label-flipping and shuffling attacks. The outputs of these meta-classifiers are then used to cancel the Byzantine encrypted updates by reweighting. Second, we propose an automated method for selecting the optimal kernel and the dimensionality hyperparameters with respect to homomorphic inference, aggregation constraints and efficiency over the CKKS cryptosystem. Finally, we demonstrate through extensive experiments the effectiveness of our approach against Byzantine participants on the $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{GTSRB}$, and $\mathsf{ACSIncome}$ benchmarks. More precisely, our SVM filtering achieves accuracies between 9090% and 9494% for identifying Byzantine updates at the cost of marginal losses in model utility and encrypted inference runtimes ranging from 66 to $24$ seconds and from 99 to $26$ seconds for an overall aggregation.
@@ 1 Introduction
    We conduct extensive experiments to demonstrate the performance of our approach against Byzantines on the $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{GTSRB}$ and $\mathsf{ACSIncome}$ datasets, along with ResNet He et al. (2015), VGG Simonyan and Zisserman (2015), custom CNN and MLP model architectures  [^1] Our implementations are available at https://github.com/Akram275/FL_with_FHE_filtering. For instance, our filtering approach achieves F1-scores between 9090% and 9494%, at the cost of marginal losses in model utility and encrypted inference runtime ranging from 33 to 77 seconds for a global aggregation (cf. Section 9).
@@ 5 Encrypted Property Filtering Framework
FL with FHE property filtering. Before training, workers initialize a threshold CKKS scheme with parameter $\tau>B_{\text{coll}}$, each holding a secret key share $\mathsf{sk}_{i}$ and the public key pk. The evaluation and public keys $(\textsf{evk},\textsf{pk})$ are sent to the server with the collaborative decryption requiring at least $\tau$ shares. The threshold CKKS setup is detailed in Appendix A.2.
When the FL training begins ($t=0$), the server broadcasts the initial plaintext model. Afterwards, workers compute local updates $\mathsf{LocalUpdate(\cdot)}$, encrypt them under pk and transmit the ciphertexts to the server. Then, the server computes the encrypted difference $\llbracket\Delta^{t}_{i}\rrbracket=\llbracket\theta^{t,s}\rrbracket-\llbracket\theta^{t,s}_{i}\rrbracket$ in which $s$ is the identifier of the layer with the highest separability (c.f., Section 6). An FHE property inference process $\mathsf{Filter}_{P}(\cdot)$ on $\llbracket\Delta^{t}_{i}\rrbracket$ results in an encrypted filter value $\llbracket P_{i}^{t}\rrbracket$ in which $P_{i}^{t}\in[0,1]$, is the filter value for worker ii. Each update is homomorphically scaled by its corresponding filter value, and the scaled updates are aggregated (via averaging) to form the encrypted non-normalized model $\llbracket\theta_{\text{scaled}}^{t+1}\rrbracket$. The server homomorphically computes $\llbracket S^{t}\rrbracket=\sum_{i=1}^{n}\frac{n_{i}}{\sum_{i=1}^{n}n_{i}}\llbracket P^{t}_{i}\rrbracket$ and its inverse $\llbracket S_{\text{inv}}^{t}\rrbracket$ to normalize the global model. Hence, producing $\llbracket\theta^{t+1}\rrbracket$, as outlined in Algorithm 4. Workers collaboratively decrypt the aggregated model using their private keys shares $\mathsf{sk}_{i}$. Concurrently, the server bootstraps the aggregated model’s ciphertext $s$ to compute the next round’s $\Delta^{t+1}_{i}$.
@@ 7.2 Optimizing the SVM Filtering
To showcase our approach, we consider four models for $\mathsf{FEMNIST}$, $\mathsf{CIFAR10}$, $\mathsf{ACSIncome}$ and $\mathsf{GTSRB}$ presented in Table III. Figure 4 depicts the enhanced separability brought by SPCA compared to PCA on backdoor update differences from the first convolution layer. Indeed, SPCA widens the separation along the first component between honest and backdoor update clusters, moving from an average distance of around 22 to more than $200$. This separation directly impacts the SVM’s decision boundary and reduces the number of support vectors, thereby improving inference runtime.
@@ 7.3 Optimal filters via grid search
Since the evaluation of the sum over all support vectors $(\mathsf{SVs})$ during inference is highly parallelizable, the factor $\mathsf{P}$ adjusts the influence of $(n_{\mathsf{sv}})$ count on the overall score. This ensures that configurations with a higher number of support vectors are not unfairly penalized when parallel processing is achievable. The weight $\mathsf{P}$ ranges from 11, indicating no parallelization capabilities, to $\frac{1}{n_{\mathsf{sv}}}$, representing maximum parallelization, in which each support vector is evaluated simultaneously, with intermediate values capturing varying parallelization capabilities.
@@ 8.2 Polynomial approximation of Heavyside
This process reduces the multiplicative complexity of the polynomial evaluation from $O(d)$ to $O(\sqrt{d})$, which leads to two key benefits. First, since multiplications are the most computationally expensive operation in CKKS (and FHE in general), reducing the number of multiplications directly enhances efficiency. Second, by lowering the overall multiplicative depth of our FHE circuit, we can use more efficient CKKS parameters, reducing the cost of all unitary FHE operations. For our Chebyshev approximation of $\mathsf{Sigmoid}$, we use $d=32$, which requires a homomorphic multiplicative depth of 88.
@@
in which $\mathsf{B}\neq 0$ denotes a new boundary value. Negative values induce less tolerance, in which higher confidence is needed to achieve $P^{t}_{j}\geq 0.5$, while positive values reduce the tolerance of the filtering process. From an FHE perspective, this is a ciphertext-plaintext addition [^3] The $\mathsf{B}$ value is encoded as a degree NN polynomial and summed directly to the second polynomial term of the RLWE ciphertext format (c.f. Appendix A)., which is computationally negligible and does not increase the ciphertext’s noise, nor consumes FHE-computational levels. This flexibility allows the approach to be tailored to the specific property. For instance, when targeting gradient-ascent attackers, incorporating even a single Byzantine update into the aggregation can severely disrupt the achieved learning progress, making it a greater concern than temporarily down-weighting an honest worker for one round.
@@ 8.3 Model Normalization with Newton-Raphson
In Algorithm 2, each worker’s update $\theta^{t}_{i}$ is scaled by $\frac{n_{i}}{\sum_{k\in[n]}n_{k}}\llbracket P^{t}_{i}\rrbracket$. This approach requires a normalization step after computing the sum of these scaled updates. Indeed, the scaling factors do not necessarily sum to 11 (unless all the property filters are equal to 11). The encrypted sum of the factors $\llbracket S^{t}\rrbracket=\sum_{i\in[n]}\left(\frac{n_{i}}{\sum_{k\in[n]}n_{k}}\right)\llbracket P^{t}_{i}\rrbracket$ is in $[0,1]$. Hence, to readjust the aggregated model’s parameters to the correct scale, a division by $S^{t}$ is required at each iteration. Since the division is not directly supported in the CKKS encryption, we approximate the inverse of $S^{t}$ using the Newton-Raphson method.
@@
Figure 8 illustrates that $k=2$ yields a sufficiently accurate inverse approximation, with an expected approximation error of around $10^{-2}$ and a computational cost of 44 FHE products. Thus, we consistently use $k=2$ for $\mathsf{EncryptedInverse}(\cdot)$ in our aggregation circuit.
@@ 9.1 Encrypted property-filtering
Table VI illustrates the practical applicability of our filtering method, showing that most FHE operations (excluding bootstrapping) are completed in $0.9$ to $\approx 15$ seconds. The SPCA projection, which processes raw, high-dimensional encrypted update vectors, is the most computationally expensive step. In contrast, the SVM hyperplane projection is much faster thanks to its lower input dimension. The Paterson-Stockmeyer optimization significantly speeds up the evaluation of the $32$-degree Chebyshev approximation of the $\mathsf{Sigmoid}$ function. As a result, the overall filter computation time remains under 99 seconds for Backdoor, gradient-ascent, and label-shuffling SVMs. Label-flipping filtering requires a quadratic kernel SVM with a larger number of support vectors and parameters ($\mathsf{P2}$), leading to an increased computation time. Although bootstrapping is costly, it is performed on the server side and can overlap with worker-side operations, thereby reducing its impact on the FL workflow. Further details are given in Appendix B.
@@ A.1 The CKKS scheme
  • $\mathsf{ParamsGen}(\lambda,L,\text{precision})$: Generates the public parameters for the CKKS scheme based on the security parameter $\lambda$, the desired multiplicative depth $L$, and the target precision for approximate arithmetic. Outputs typically include: the polynomial ring degree NN (a power of 2, determining the vector length), the chain of ciphertext coefficient moduli $Q=q_{0}\cdot q_{1}\cdots q_{L}$ (where $Q_{l}=q_{0}\cdots q_{l}$ is the modulus at level $l$), an auxiliary modulus $P$ used for key switching (relinearization, rotation), the scaling factor $\Delta$ (often a power of 2, related to precision), and specifications for the secret key distribution $S$ and error distribution $\chi$. These parameters must ensure $\lambda$-bit security while supporting $L$ levels of multiplication and the desired precision.
@@
  • $\mathsf{RelinKeyGen}(\textsf{sk})$: Generates the relinearization key $\mathsf{rlk}$, which is necessary to reduce the size of ciphertexts after multiplication. It is effectively an encryption of $s^{2}$ under the secret key sk, typically constructed using the extended modulus PQPQ. A common form involves sampling $a^{\prime}\leftarrow R_{PQ}$, $e^{\prime}\leftarrow\chi$, yielding $\mathsf{rlk}=(b^{\prime},a^{\prime})=([-a^{\prime}s+e^{\prime}+Ps^{2}]_{PQ},a^{\prime})$.
  • $\mathsf{RotKeyGen}(\textsf{sk},k)$: Generates a rotation key $\mathsf{rotk}_{k}$ required for homomorphically rotating the encrypted vector by $k$ positions (where $k$ corresponds to a specific Galois element). This key effectively encrypts the modified secret key $s(X^{k})$ under the original secret key sk, typically using the modulus PQPQ. Keys must be generated for all desired rotation indices $k$.
@@ Appendix B Offline amortized cost of Bootstrapping
During the collaborative decryption phase, workers generate $t$-out-$t$ additive keys by (at least) $t$ active parties, using their respective $n$ shares of the initial $n$ additive secret keys (c.f. equation 4). Apart from revealing the active parties, the process of computing the $t$ partial decryptions does not require data exchanged other than what has already been exchanged during the setup. The last step is broadcasting the respective partial decryption by each of the $t$ active parties. A partially decrypted message consists of a degree NN polynomial with coefficients in $\mathbb{Z}_{Q}$. Thus, the size of data exchanged (for minimal $t$ active parties) is $t(t-1)N\log_{2}(Q)$ bits for a single ciphertext. Encrypting a model update requires $\left\lceil\frac{\mathsf{dim}(\theta)}{N/2}\right\rceil$ ciphertexts, which are 3 to 15 ciphertexts for our models (see table III). Table VII illustrates the size of the exchanged data during the collaborative decryption and the runtime necessary to generate the shares for several values of $\log(N)$, and different numbers of active parties ($t$).